Not wanting to be embarrassed by revealing her age on her birthday, Grandma instead told the grandkids that if they multiplied her age in 5 years by 5, then multiplied her age in 6 years by 6, and added the two totals together, they would get a number that is 12 times her current age. When she saw the large wax forest fire atop her birthday cake , Grandma knew they had figured it out. How old is she?
Answer by mananth(16946)

Not wanting to be embarrassed by revealing her age on her birthday, Grandma instead told the grandkids that if they multiplied her age in 5 years by 5, then multiplied her age in 6 years by 6, and added the two totals together, they would get a number that is 12 times her current age. When she saw the large wax forest fire atop her birthday cake , Grandma knew they had figured it out. How old is she?
let her age be x now
after 5 years
(x+5)*5
after 6 years
(x+6)*6

5(x+5) +6(x+6) = 12x
5x+25+6x+36=12x
11x+61=12x
12x-11x=61
x=61